On a hot summer day in the state of Washington while kayaking, I saw several swimmers jump from a railroad bridge into the Snohomish River below. The swimmers stepped off the bridge, and I estimated that they hit the water 2.00 s later. You may want to review (Pages 44 - 49)
Answer #1

Given

t = 2 seconds

Here, we need to use kinematics,

(a) h = 1/2at2

here a = g = 9.8 m/s2

h = 1/2 * 9.8 * 22

h = 19.6 m

------------------------------------------------------------------

(b) v2 - u2 = 2gh

assuming, the swimmer jumps from rest, u = 0

so,

v = sqrt (2gh)

v = sqrt ( 2 * 9.8 * 19.6)

v = 19.6 m/s